OpenStudy (anonymous):

Which of the following are AND "point-slope" AND equations AND of AND the AND line AND going AND through (4 6) and (1 2)

OpenStudy (texaschic101):

first lets find the slope.. slope (m) = (y2 - y1)/(x2-x1) set 1 (4,6) set 2 (1,2) m = (2 - 6) / (1 - 4) m = -4/-3 m = 4/3 y = mx + b using slope 4/3 and either of the points...I am using (4,6) 6 = 4/3(4) + b 6 = 16/3 + b 6 - 16/3 = b 18/3 - 16/3 = b 2/3 = b your equation is y = 4/3x + 2/3

OpenStudy (texaschic101):

oh...sorry...you want point slope form... y - y1 = m(x - x1) using slope 4/3 and either of the points (4,6) y - 6 = 4/3(x - 4) < -- point slope form using point (1,2) y - y1 = m(x - x1) y - 2 = 4/3(x - 1) < -- point slope form